# Robustness-Aware 3D Object Detection in Autonomous Driving: A Review and Outlook

### v1

@@ IV-E1 Accuracy
Referencing Section III discussion on camera-based methods, to investigate the core factors influencing the performance of LiDAR-based methods similarly, we selected representative and cutting-edge methods from each category for a comparative performance analysis, as shown in Fig 10. The current scenario indicates that the performance of view-based methods is significantly lower than the other three categories. view-based approaches aim to transform point clouds into pseudo-images, then process data using 2D detectors. Although this is favorable for inference speed, it comes at the cost of sacrificing some 3D spatial information. As discussed in the section III, camera-based detection methods that can effectively recover 3D space often exhibit superior detection performance. For example, the detection performance of binocular methods surpasses that of monocular methods, and methods based on the BEV paradigm outperform others. Therefore, the effective representation of 3D spatial information is a primary factor impacting the performance of LiDAR-based methods.
In the development of LiDAR-based detection methods, initially, point-based and PV-based approaches outperformed voxel-based methods. However, over time, methods represented by Voxel RCNN [113], utilizing ROI pool modules capable of effectively aggregating fine-grained information, have brought comparable or superior performance to voxel-based approaches. The distinction of the Voxel-RCNN approach from earlier voxel-based methods lies in its ROI pooling module, which effectively aggregates local fine-grained information, addressing the issue of loss of detailed 3D spatial information in the voxelization process. As shown in Fig. 6(c) and (d), the comparison between single-stage [109, 118] and two-stage [113, 127] LiDAR methods indicates that the latter’s use of fine-grained data significantly enhances accuracy. This factor also contributed to the early advantage of point-based and PV-based methods over voxel-based approaches. However, recent voxel-based methods, focusing on extracting and utilizing detailed point cloud information, have narrowed this gap in precision. This phenomenon underscores the importance of fine-grained point cloud information in improving detection accuracy.
Additionally, it is important to note the accuracy differences between transformer-based methods and those based on CNNs or PointNet, as shown in Fig. 10. Due to both global and local self-attention mechanisms that capture contextual information from a large spatial range, transformers have an advantage in modeling long-range dependencies. The receptive field of 3D sparse CNNs is constrained by the size of the convolutional kernel and the spatial discontinuities of sparse features, leading to information truncation issues and thus inferior performance compared to transformers. Recent related works [130, 129] provide a good demonstration that when the receptive field of 3D convolutions is expanded, methods based on 3D sparse convolutions can achieve the performance of transformer-based methods. Therefore, effective global relationship modeling and avoiding information truncation are other key factors in enhancing the accuracy of LiDAR-based detection.
@@ IV-E2 Latency
The section III highlights latency’s importance in autonomous driving safety and user experience. While camera-based methods tend to outperform LiDAR-based methods in terms of inference speed, the latter still maintain a competitive edge due to their accurate 3D perception. We conducted tests using an A100 graphics card to measure the FPS of significant LiDAR-based approaches, and evaluated their performance using the original research’s AP and NDS metrics. As shown in Fig. 6, it indicates that view-based methods excel in model latency due to the reduction in point cloud dimensions and the efficiency of 2D CNNs. Voxel-based methods achieve exceptional inference speed due to the use of structured voxel data and well-optimized 3D sparse convolutions. However, point-based methods face challenges in applying efficient operators during data preprocessing and feature extraction stages due to the irregular representation of point clouds. Point-GNN [110] is an extreme example of this, with model latency nearly ten times that of contemporary voxel algorithms.
To conclude, optimizations improving detection performance typically compromise inference speed in autonomous driving models. Achieving a balance between accuracy and speed is an evolving challenge in this field. Future studies should prioritize the simultaneous improvement of accuracy, FPS and latency reduction in order to meet the urgent requirements of real-time response and safety in autonomous driving.
@@ IV-E3 Robustness
Previous comprehensive reviews have not focused significantly on the topic of robustness. Presently, research works[253, 254, 388, 287, 288, 389, 390] like RoboBEV[253], Robo3D[288] on 3D object detection incorporate considerations of robustness, exemplified by factors such as sensor misses. Robo-LiDAR [286] represents the first comprehensive exploration solely dedicated to the robustness of LiDAR-based methods. In a manner akin to BR3D [254], this method evaluates robustness by integrating disturbances into datasets pertinent to 3D object detection, such as KITTI [255]. The method involves proposing a variety of noise types and 25 typical degradations associated with object and scene-level natural weather conditions, noise interferences, density variations, and object transformations. In this section, we will combine the work of Ref. [254] and Robo-LiDAR [286] with the aim of systematically analyzing the robustness of LiDAR-based methods. As shown in the Table IV, generally, LiDAR-based methods exhibit higher robustness to noise compared to camera-based methods. It is important to note the performance differences between LiDAR-based methods and multimodal methods in noisy environments. In multimodal methods [227, 232, 230, 221, 228, 247, 237], the complementary interplay of data types becomes evident when disturbances are limited to LiDAR sensor data. In such scenarios, image data can partially mitigate the impact on point cloud integrity, consequently elevating the performance of fusion methods above that of methods relying solely on LiDAR. when disturbances affect both image and point cloud data concurrently, the efficacy of most multimodal methods significantly diminishes.
@@
As shown in the Table VI, under various noise conditions, LiDAR-based methods experience varying degrees of accuracy decline, with the most significant reduction observed in extreme weather noise scenarios. These results indicate an urgent need in the field of autonomous driving to address the robustness issue of point cloud detectors. For most types of corruptions, voxel-based methods generally exhibit greater robustness than point-based methods. A plausible explanation is that voxelization, through the spatial quantization of a group of adjacent points, mitigates the local randomness and spatial information disruption caused by noise and density degradation. Specifically, for severe corruptions (e.g., shear, FFD in the Transformation), the point-voxel-based method [371] exhibits greater robustness. PointRCNN [111] does not show the highest robustness against any form of corruptions, highlighting potential limitations inherent in point-based methods.
Compared to single-stage detectors [109, 143, 127], two-stage detectors [125, 111, 371, 142, 127] exhibit poorer robustness to common degradations, as indicated by lower scores. One possible cause is that corrupted data could affect the proposal generation of stage 1, and the low-quality proposals significantly affect the BBox regression of stage 2. In summary, single-stage detectors demonstrate greater robustness against scene-level corruptions and object-level corruptions and density disruptions, while two-stage detectors are mainly more robust against weather and scene-level density degradations. As for Transformation corruptions, one-stage detectors present better robustness on FFD, rotation, translation and two-stage detectors work better under corruptions of shear, scale.
| Corruption | Point-voxel | Point | Voxel | Average | |||||||
| PVRCNN-voxel | PointRCNN | SECOND | BtcDet | VoTr-SSD | VoTr-TSD | SE-SSD | CenterPoint | ||||
| Scene-level | Weather | rain | 25.11 | 23.31 | 21.81 | 31.07 | 28.17 | 26.77 | 29.51 | 25.83 | 26.45 |
| snow | 44.23 | 37.74 | 34.84 | 54.07 | 54.10 | 52.18 | 49.19 | 38.74 | 45.64 | ||
| fog | 1.59 | 3.52 | 1.60 | 1.81 | 1.77 | 2.02 | 1.59 | 1.11 | 1.88 | ||
| Noise | uniform_rad | 10.19 | 8.32 | 9.51 | 9.13 | 3.79 | 4.11 | 9.34 | 8.15 | 7.82 | |
| gaussian_rad | 13.02 | 9.98 | 12.13 | 10.83 | 4.84 | 5.18 | 11.02 | 10.17 | 9.65 | ||
| impulse_rad | 2.20 | 3.86 | 2.23 | 2.50 | 2.25 | 3.57 | 1.18 | 1.8 | 2.46 | ||
| background | 2.93 | 6.49 | 2.41 | 1.82 | 4.59 | 3.6 | 2.14 | 1.86 | 2.46 | ||
| upsample | 0.81 | 1.84 | 0.31 | 0.95 | 0.37 | 0.71 | 0.55 | 0.46 | 0.75 | ||
| Density | cutout | 3.75 | 3.97 | 4.27 | 3.99 | 4.51 | 3.59 | 4.26 | 4.11 | 4.0 | |
| local_dec | 14.04 | - | 13.88 | 14.55 | 14.44 | 12.50 | 17.01 | 14.64 | 14.44 | ||
| local_inc | 1.40 | 3.34 | 1.33 | 2.20 | 1.66 | 1.69 | 0.90 | 0.95 | 1.68 | ||
| beam_del | 0.58 | 0.79 | 0.73 | 0.88 | 0.80 | 0.53 | 1.07 | 0.47 | 0.73 | ||
| layer_del | 2.94 | 3.46 | 3.10 | 3.39 | 3.29 | 3.16 | 3.37 | 2.67 | 3.17 | ||
| Object-level | Noise | uniform | 15.44 | 12.95 | 9.48 | 12.6 | 2.76 | 4.81 | 6.99 | 6.51 | 8.94 |
| gaussian | 20.48 | 17.62 | 12.98 | 17.05 | 4.72 | 7.46 | 9.56 | 9.49 | 12.42 | ||
| impulse | 3.3 | 4.7 | 2.53 | 4.07 | 2.88 | 4.29 | 2.2 | 2.11 | 3.26 | ||
| upsample | 1.12 | 1.95 | 0.67 | 1.33 | 0.08 | 0.4 | 0.22 | 0.16 | 0.74 | ||
| Density | cutout | 15.81 | 15.62 | 14.99 | 15.62 | 15.07 | 16.09 | 16.51 | 14.06 | 15.47 | |
| local_dec | 14.38 | 14.16 | 13.23 | 14.26 | 12.66 | 14.41 | 15.08 | 12.52 | 13.84 | ||
| local_inc | 13.93 | 14.19 | 13.74 | 13.56 | 11.34 | 13.05 | 11.03 | 11.64 | 12.81 | ||
| Transformation | shear | 37.27 | 40.96 | 40.35 | 41.37 | 39.52 | 37.85 | 40.35 | 40.0 | 39.71 | |
| FFD | 32.42 | 38.88 | 33.15 | 36.77 | 33.14 | 34.26 | 37.96 | 32.86 | 34.93 | ||
| rotation | 0.60 | 0.47 | 0.31 | 0.97 | 0.39 | 0.75 | 0.27 | 0.38 | 0.52 | ||
| scale | 5.78 | 8.13 | 6.96 | 5.81 | 8.53 | 6.50 | 6.53 | 7.50 | 6.97 | ||
| translation | 3.82 | 3.03 | 3.24 | 4.58 | 4.88 | 5.34 | 1.37 | 3.91 | 3.77 | ||
| mCE | 11.49 | 11.64 | 10.39 | 12.21 | 10.42 | 10.60 | 11.17 | 10.09 | 11.01 | ||
| Corruption | One-stage | Two-Stage | ||
| Sence-level | Weather | rain | 26.50 | 26.42 |
| snow | 46.04 | 45.39 | ||
| fog | 2.01 | - | ||
| Noise | uniform_rad | 7.55 | 7.98 | |
| gaussian | 9.33 | 9.84 | ||
| impulse_rad | 1.30 | 1.02 | ||
| background | 3.05 | 3.38 | ||
| upsample | 0.41 | 0.95 | ||
| Density | cutout | 4.35 | 3.88 | |
| local_dec | 15.12 | 13.93 | ||
| local_inc | 1.30 | 1.92 | ||
| beam_del | 0.87 | 0.65 | ||
| layer_del | 3.25 | 3.12 | ||
| Object-level | Noise | uniform | 6.41 | 10.46 |
| gaussian | 9.09 | 14.42 | ||
| impulse | 2.54 | 3.69 | ||
| upsample | 0.32 | 0.99 | ||
| Density | cutout | 15.52 | 15.44 | |
| local_dec | 13.66 | 13.95 | ||
| local_inc | 12.04 | 13.27 | ||
| Transformation | shear | 40.07 | 39.49 | |
| FFD | 34.75 | 35.04 | ||
| rotation | 0.32 | 0.63 | ||
| scale | 7.34 | 6.74 | ||
| translation | 3.16 | 4.14 | ||
| mCE | 10.66 | 11.21 | ||
V multimodal 3D Object Detection
@@
multimodal 3D object detection refers to the technique of using data features from different sensors and integrating these features to achieve complementarity, thus enabling the detection of 3D objects. As shown in Fig. 11, they particularly emphasizes the combination of image data and point cloud data. Image data is rich in semantic features, such as color and texture, but often lacks depth information. In contrast, point cloud data provides depth information and geometric structure, which is crucial for accurately perceiving and interpreting the 3D characteristics of a scene. Since a single type of sensor cannot fully and accurately perceive the 3D environment, multimodal 3D object detection acquires features with rich semantic information by fusing various types of data.
@@
In the field of autonomous driving, there are a variety of fusion methods for multimodal 3D object detection. Previous reviews [1, 344, 345, 346, 347, 242, 2, 348, 349, 350, 269, 252, 270, 271, 272, 273, 274, 275] have mostly classified these methods based on different stages of fusion (early, middle, late), but this classification is overly simplistic and does not fully consider the special requirements of autonomous driving. Given the fundamental differences between the two heterogeneous modalities of point clouds and images, the alignment step in multimodal fusion is particularly critical. It ensures the consistency and accuracy of information from different sensors and data sources during the fusion process. In autonomous driving, the key to achieving feature alignment lies in whether to use a calibration matrix (also known as a projection matrix). However, the inherent error of the calibration matrix, being a type of prior knowledge, poses a challenge. Some works like [437, 236] avoid using the projection matrix and reduce projection errors by adopting learning methods.
@@
Therefore, based on different methods of feature alignment, we can categorize multimodal 3D object detection methods into: (1) projection-based for feature alignment, (2) model-based for feature alignment. This taxonomy is more detailed and scientific, better reflecting the characteristics and progress of multimodal 3D object detection methods in the field of autonomous driving.
@@ V-A Projection-based 3D object detection
Projection-based 3D object detection refers to the use of projection matrix during the feature fusion stage to achieve the integration of point cloud and image features. It is important to clarify that the focus here is on projection during the feature fusion period, rather than those in the fusion stage, which includes projections needed for processes such as data augmentation. As shown in Fig. We have developed a more detailed classification of Projection-based 3D object detection based on the different types of projection used in the fusion stage, including Point-Projection-based [222, 383, 438, 439, 440, 441, 442, 226, 227, 376, 443], Feature-Projection-based [444, 250, 232, 445, 446, 447, 448, 449, 377], Auto-Projection-based [369, 106, 233, 234, 384, 378, 243, 247, 450],[369] and Decision-Projection-based methods [451, 452, 453, 380, 381, 196, 454, 382, 107].
@@ V-A1 Point-Projection-based 3D object detection
Point-Projection-based 3D object detection methods [222, 383, 438, 439, 440, 441, 442, 443] involve a process of projecting image features onto raw point clouds to enhance the representational capability of the original point cloud data. The initial step in these methods is to establish a strong correlation between LiDAR points and image pixels, which is achieved using calibration matrices. Following this, the point cloud features are enhanced by augmenting them with additional data. This augmentation takes two forms: either through the incorporation of segmentation scores [222, 438, 456] or by using CNN features [376, 443, 226, 227, 383] from the correlated pixels. PointPainting [222] and PointAugmenting [383] represent advancements in multi-model 3D object detection methods by enhancing the traditional cut-and-paste augmentation. These techniques aim to seamlessly integrate data from different domains, such as point clouds and 2D imagery, while carefully managing potential overlaps or collisions between objects in both domains. PointPainting enhances LiDAR points by appending segmentation scores. However, it has limitations in effectively capturing the color and texture details present in images. To address these shortcomings, more sophisticated approaches like FusionPainting [456] have been developed, following a similar paradigm. MVP [443] builds upon the concept of PointPainting[222]. It initially utilizes image instance segmentation and establishes an alignment between the segmentation masks and the point cloud using a projection matrix. The key distinction of MVP lies in its approach to sampling: it randomly selects pixels within each range, ensuring consistency with the points in the point cloud. These selected pixels are then linked to their nearest neighbors in the point cloud. The depth value of the LiDAR point in this linkage is assigned as the depth of the corresponding pixel. Subsequently, these points are projected back to the LiDAR coordinate system, resulting in the generation of virtual LiDAR points.

### v2

@@ IV-E1 Accuracy
Referring to Section III on Camera-only methods, we investigated the core factors influencing LiDAR-only methods. Representative methods from each category underwent comparative performance analysis on the KITTI [133] and nuScenes [134] datasets, as shown in Fig 3 (c, d). The current scenario indicates that the latest view-based methods exhibit lower performance compared to other categories. View-based approaches transform point clouds into pseudo-images for processing with 2D detectors, which favors inference speed but sacrifices 3D spatial information. Therefore, an effective representation of 3D spatial information is pivotal for LiDAR-only methods. Initially, point-based and PV-based methods outperformed voxel-based approaches in LiDAR-only detection. Over time, methods like Voxel RCNN [47], which utilize ROI pool modules for fine-grained information aggregation, have achieved comparable or superior performance. Voxel RCNN’s ROI pooling module effectively addresses the loss of detailed 3D spatial information resulting from voxelization.
@@ IV-E2 Latency
Section III highlights latency’s importance in autonomous driving safety and user experience. While Camera-only methods tend to outperform LiDAR-only methods in terms of inference speed, the latter still maintain a competitive edge due to their accurate 3D perception. We conducted tests using an A100 graphics card to measure the FPS of significant LiDAR-only approaches, and evaluated their performance using the original research’s AP and NDS metrics. As shown in Table V, it indicates that view-based methods excel in model latency due to the reduction in point cloud dimensions and the efficiency of 2D CNNs. Voxel-based methods achieve exceptional inference speed due to the use of structured voxel data and well-optimized 3D sparse convolutions. However, point-based methods face challenges in applying efficient operators during data preprocessing and feature extraction stages due to the irregular representation of point clouds. Point-GNN [44] is an extreme example of this, with model latency nearly several times that of contemporary voxel algorithms. Transformer-based methods [67] face significant challenges in real-time inference. The current research trend in transformer-based methods is the development of efficient attention operators, like [55, 64, 371]. Moreover, for PV-based methods, the construction of point-to-voxel or voxel-topoint relationships, along with the feature fusion of voxels and points, incurs additional computational overhead. To conclude, common accuracy optimization strategies, such as two-stage optimization or attention mechanisms, typically compromise inference speed in autonomous driving models. Achieving a balance between accuracy and speed is an evolving challenge in this field. Future studies should prioritize the simultaneous improvement of accuracy, as well as the reduction of FPS (frames per second) and latency, in order to meet the urgent requirements of real-time response and safety in autonomous driving.
@@ IV-E3 Robustness
Previous comprehensive reviews have not focused significantly on the topic of robustness. Presently, research works[124, 125, 390, 147, 148, 391, 392] like RoboBEV[124], Robo3D[148] on 3D object detection incorporate considerations of robustness, exemplified by factors such as sensor misses. Robo-LiDAR [146] represents the first comprehensive exploration solely dedicated to the robustness of LiDAR-only methods. In a manner akin to BR3D [125], this method evaluates robustness by integrating disturbances into datasets pertinent to 3D object detection, such as KITTI [133]. The method involves proposing a variety of noise types and 25 typical degradations associated with object and scene-level natural weather conditions, noise interferences, density variations, and object transformations. In this section, we will combine the work of Ref. [125] and Robo-LiDAR [146] with the aim of systematically analyzing the robustness of LiDAR-only methods. As shown in the Table VII, generally, LiDAR-only methods exhibit higher robustness to noise compared to Camera-only methods. In Multi-modal methods [105, 110, 108, 99, 106, 121, 116], the complementary interplay of data types becomes evident when disturbances are limited to LiDAR sensor data. In such scenarios, image data can partially mitigate the impact on point cloud integrity, consequently elevating the performance of fusion methods above that of methods relying solely on LiDAR. when disturbances affect both image and point cloud data concurrently, the efficacy of most Multi-modal methods significantly diminishes. It is worth noting that DTS [416] and Bi3D [417] enhance model robustness through domain adaptation methods.
@@
As shown in Table IX, under various noise conditions, LiDAR-only methods experience varying degrees of accuracy decline, with the most significant reduction observed in extreme weather noise scenarios. These results indicate an urgent need in the field of autonomous driving to address the robustness issue of point cloud detectors. For most types of corruptions, voxel-based methods generally exhibit greater robustness than point-based methods, as shown in Table VII, VIII, IX. A plausible explanation is that voxelization, through the spatial quantization of a group of adjacent points, mitigates the local randomness and spatial information disruption caused by noise and density degradation. Specifically, for severe corruptions (e.g., shear, FFD in the transformation), the point-voxel-based method [308] exhibits greater robustness. PointRCNN [45] does not show the highest robustness against any form of corruption, highlighting potential limitations inherent in point-based methods. In conclusion, future works should explore robustness optimization from the perspectives of data representation and model architecture. The above analysis aims to offer valuable insights for future work related to robustness.
| Corruption | PV | Point | Voxel | Avg. | ||||
| PV-RCNN | PointRCNN | SECOND | SE-SSD | CenterPoint | ||||
| Scene-level | Weather | rain | 25.11 | 23.31 | 21.81 | 29.51 | 25.83 | 26.45 |
| snow | 44.23 | 37.74 | 34.84 | 49.19 | 38.74 | 45.64 | ||
| fog | 1.59 | 3.52 | 1.60 | 1.59 | 1.11 | 1.88 | ||
| Noise | uniform_rad | 10.19 | 8.32 | 9.51 | 9.34 | 8.15 | 7.82 | |
| gaussian_rad | 13.02 | 9.98 | 12.13 | 11.02 | 10.17 | 9.65 | ||
| impulse_rad | 2.20 | 3.86 | 2.23 | 1.18 | 1.8 | 2.46 | ||
| background | 2.93 | 6.49 | 2.41 | 2.14 | 1.86 | 2.46 | ||
| upsample | 0.81 | 1.84 | 0.31 | 0.55 | 0.46 | 0.75 | ||
| Density | cutout | 3.75 | 3.97 | 4.27 | 4.26 | 4.11 | 4.0 | |
| local_dec | 14.04 | - | 13.88 | 17.01 | 14.64 | 14.44 | ||
| local_inc | 1.40 | 3.34 | 1.33 | 0.90 | 0.95 | 1.68 | ||
| beam_del | 0.58 | 0.79 | 0.73 | 1.07 | 0.47 | 0.73 | ||
| layer_del | 2.94 | 3.46 | 3.10 | 3.37 | 2.67 | 3.17 | ||
| Object-level | Noise | uniform | 15.44 | 12.95 | 9.48 | 6.99 | 6.51 | 8.94 |
| gaussian | 20.48 | 17.62 | 12.98 | 9.56 | 9.49 | 12.42 | ||
| impulse | 3.3 | 4.7 | 2.53 | 2.2 | 2.11 | 3.26 | ||
| upsample | 1.12 | 1.95 | 0.67 | 0.22 | 0.16 | 0.74 | ||
| Density | cutout | 15.81 | 15.62 | 14.99 | 16.51 | 14.06 | 15.47 | |
| local_dec | 14.38 | 14.16 | 13.23 | 15.08 | 12.52 | 13.84 | ||
| local_inc | 13.93 | 14.19 | 13.74 | 11.03 | 11.64 | 12.81 | ||
| T.F. | shear | 37.27 | 40.96 | 40.35 | 40.35 | 40.0 | 39.71 | |
| FFD | 32.42 | 38.88 | 33.15 | 37.96 | 32.86 | 34.93 | ||
| rotation | 0.60 | 0.47 | 0.31 | 0.27 | 0.38 | 0.52 | ||
| scale | 5.78 | 8.13 | 6.96 | 6.53 | 7.50 | 6.97 | ||
| translation | 3.82 | 3.03 | 3.24 | 1.37 | 3.91 | 3.77 | ||
| mCE | 11.49 | 11.64 | 10.60 | 11.17 | 10.09 | 11.01 | ||
V Multi-modal 3D Object Detection
@@
Multi-modal 3D object detection refers to the technique of using data features from different sensors and integrating these features to achieve complementarity, thus enabling the detection of 3D objects. As shown in Fig. 6, the approach particularly emphasizes the combination of image data and point cloud data. Image data is rich in semantic features, such as color and texture, but often lacks depth information. In contrast, point cloud data provides depth information and geometric structure, which is crucial for accurately perceiving and interpreting the 3D characteristics of a scene. Since a single type of sensor cannot fully and accurately perceive the 3D environment, multi-modal 3D object detection acquires features with rich semantic information by fusing various types of data.
@@
In the field of autonomous driving, there are a variety of fusion methods for multi-modal 3D object detection. Previous reviews [4, 126, 123, 130, 131] have mostly classified these methods based on different stages of fusion (early, middle, late), but this classification is overly simplistic and does not fully consider the special requirements of autonomous driving. Given the fundamental differences between the two heterogeneous modalities of point clouds and images, the alignment step in multi-modal fusion is particularly critical. It ensures the consistency and accuracy of information from different sensors and data sources during the fusion process. In autonomous driving, the key to achieving feature alignment lies in whether to use a calibration matrix (also known as a projection matrix). However, the inherent error of the calibration matrix, being a type of prior knowledge, poses a challenge. Some works, like [339, 115], avoid using the projection matrix and reduce projection errors by adopting learning methods.
@@
Therefore, based on different methods of feature alignment, we can categorize multi-modal 3D object detection methods into two types: (1) projection-based for feature alignment, and (2) model-based for feature alignment. This taxonomy is more detailed and scientific, better reflecting the characteristics and progress of multi-modal 3D object detection methods in the field of autonomous driving.
@@ V-A Projection-based 3D object detection
Projection-based 3D object detection refers to the use of projection matrices during the feature fusion stage to achieve the integration of point cloud and image features. It is important to clarify that the focus here is on projection during the feature fusion period, rather than projections in other stages of the fusion process, which include projections needed for processes such as data augmentation. As shown in Fig.7, we have developed a more detailed classification of projection-based 3D object detection based on the different types of projection used in the fusion stage, including Point-Projection-based [100, 319, 320, 322, 317, 318, 418, 104, 105, 316, 321], Feature-Projection-based [325, 122, 110, 419, 327, 328, 329, 420, 326], Auto-Projection-based [296, 41, 111, 112, 332, 382, 120, 121, 331], and Decision-Projection-based methods [333, 334, 335, 383, 336, 96, 337, 338, 113].
@@ V-A1 Point-Projection-based 3D object detection
Point-Projection-based 3D object detection methods [100, 319, 320, 322, 317, 318, 418, 321] involve projecting image features onto raw point clouds to enhance the representational capability of the original point cloud data, as shown in Fig. 7 (a). The initial step in these methods is to establish a strong correlation between LiDAR points and image pixels, which is achieved using calibration matrices. Following this, the point cloud features are enhanced by augmenting them with additional data. This augmentation takes two forms: either through the incorporation of segmentation scores [100, 320, 323] or by using CNN features [316, 321, 104, 105, 319] from the correlated pixels. PointPainting [100] and PointAugmenting [319] represent advancements in multi-modal 3D object detection methods by enhancing the traditional cut-and-paste augmentation. These techniques aim to seamlessly integrate data from different domains, such as point clouds and 2D imagery, while carefully managing potential overlaps or collisions between objects in both domains. PointPainting enhances LiDAR points by appending segmentation scores. However, it has limitations in effectively capturing the color and texture details present in images. To address these shortcomings, more sophisticated approaches like FusionPainting [320] have been developed, following a similar paradigm. MVP [321] builds upon the concept of PointPainting [100]. It initially utilizes image instance segmentation and establishes an alignment between the segmentation masks and the point cloud using a projection matrix. The key distinction of MVP lies in its approach to sampling: it randomly selects pixels within each range, ensuring consistency with the points in the point cloud. These selected pixels are then linked to their nearest neighbors in the point cloud. The depth value of the LiDAR point in this linkage is assigned as the depth of the corresponding pixel. Subsequently, these points are projected back to the LiDAR coordinate system, resulting in the generation of virtual LiDAR points.